# Hug Reports: Supporting Expression of Appreciation between Users and Contributors of Open Source Software Packages

### v1

@@ 4.3.3. Procedure and Participants
To understand contributors’ reactions to the notifications, the email also included an invitation to participate in our study with a link to a survey. The survey included questions on demographics including age, gender, programming proficiency, and their tenure in the project for which they were thanked. It then asked contributors how they felt on three single-item scales adapted from prior work (Kumar and Epley, 2018): (1) a scale ranging from -5 (much more negative than normal) to 5 (much more positive than normal), with the midpoint of 0 labeled no different than normal; (2) a scale ranging from -5 (not at all surprised) to 5 (extremely surprised); and (3) a scale ranging from -5 (not at all awkward) to 5 (extremely awkward). We also included an open-ended question asking contributors to further describe how the notification made them feel. The survey included three more open ended questions: (1) Prior to this, how often did you receive messages of thanks from users?; (2) What else would you have liked to know about the senders of these thanks?; and (3) Do you have any feedback for us that you would like to share? Finally, the survey invited contributors to participate in an interview. There was no compensation for participating in the survey and all questions on the survey were indicated as optional.
@@ 4.4. Analysis
We used the usage data collected from the extension deployment as one source of data to address RQ2-U. From this, we created summary statistics (counts) of the interactions that took place. We also used an affinity diagramming approach to group the personal notes into categories, based on their content. We used contributors’ responses to the survey, as one source of data to address RQ1-C and RQ2-C. We created summary statistics (counts) for responses to quantitative questions, and used an affinity diagramming approach to analyze the responses to open-ended questions. In this way, we use these two sources of data to derive descriptive summaries of the behaviors of users and reactions of contributors in our study.
We used the interviews with users and contributors to address RQ1 (RQ1-U1, RQ1-U2, RQ1-C), RQ2 (RQ2-U, RQ2-C), and RQ3. To do this, we conducted a reflexive thematic analysis (Braun and Clarke, 2006). Two of the authors independently performed a line-by-line open coding of transcripts from the first 7 (out of the 14) user interviews and the first 5 (out of the 10) contributor interviews. Codes generated in this phase were in part inductive, driven by the data, and in part guided by our original research questions– we remained open to capturing observations that emerged through the data while also looking out for observations that related to our main guiding questions. All authors met to discuss the analysis and iteratively refined the codes, following which the first author applied the refined set of codes to the remaining transcripts looking at whether participant experiences fit into our existing categories. Finally, all authors discussed the analysis to iteratively refine and solidify the themes, and group similar themes together. Themes were generated at a semantic level, reflecting what participants explicitly said (Braun and Clarke, 2006).
@@ 5. Findings
In this section, we begin with findings from the descriptive analysis of usage data and contributors’ responses to the questionnaire and survey (5.1). In it, we provide a descriptive summary of the extension’s use and contributors’ reactions, addressing RQ2-U, RQ1-C, and RQ2-C. Across the next subsections, we present main themes from our analysis of the interviews. In section 5.2, we first discuss the extent to which key decisions of the extension encouraged meaningful appreciation (RQ1). We then present findings around when users expressed appreciation (5.3). These findings speak to how users expressed appreciation (RQ2-U). Then, we discuss the two meanings the that appreciation took on (5.4): (1) as a measure of utility, where the volume of thanks, and what is was directed at, were interpreted as an signal of the software’s utility, and (2) as an act of expressive communication that intended to convey a user’s gratitude. Users’ interactions and contributors’ interpretations depended on which of the two meanings they prioritized in a given context (5.4). These findings speak to RQ2-U and RQ2-C. Then, we discuss how contributors’ reactions we influenced by how much they felt they had contributed to the object that was being appreciated (5.5), further addressing RQ2-C. Finally, we present participants’ ideas (5.6), addressing RQ3.
@@ 5.1.1. Patterns of use (RQ2-U)
Figure 5. Summary of contributors’ responses to the survey. We include some contributors’ responses to the open ended questions in the survey to provide further context for some responses. (A) Most contributors reacted positively to the notification, and only 1 participant reacted negatively (whose survey response suggests it was because they found the thanks and personal note too generic). (B) Most contributors were surprised to receive the notification, however, few contributors who belonged to large projects were not. (C) Most contributors felt no or low levels of awkwardness being appreciated. One contributor, who felt ‘extremely awkward’ noted that receiving a ‘hug’ from a stranger felt awkward.
5.1.2. Contributors’ responses (RQ1-C, RQ2-C)
@@
Overall, contributors reacted positively to receiving appreciation. Contributors hoped to know more about why they were being thanked, and would have liked more personalized messages. Some contributors didn’t feel like they deserved the appreciation if they hadn’t substantially contributed to the package for which they were thanked. Figure 5 summarizes the responses were received to the survey, from 26 contributors. 14 contributors reported that receiving the notification made them feel more positive than normal (Figure 5 (A)), 11 reported feeling no different than normal, and only 1 participant reacted negatively (whose survey response suggests it was because they found the thanks and personal note too generic). Reponding to how often they currently received appreciation, most contributors reported it was rare, some suggesting they received “a few per year” (C15), such as in “GitHub Issues once every couple months” (C13) or “once a feature request is closed, primarily with emojis, and it is extremely rare to get a text message” (C23). Participants further described how the notification made them feel. One mentioned how “receiving the hug report was a wonderful surprise, and brightened [their] day” (C3), while another mentioned they “felt happy and feel motivated to do more” (C1). One contributor explicitly mentioned how current channels for appreciation are few: “receiving thanks for open-source work felt nice and very new since I experienced it very rarely so far. Part of why it is so rare may be that there are no convenient and established ways to do so” (C20). However, few contributors noted how there was “nothing personalized about the report” (C9). Without a more specific message, C18 mentioned: “although I do appreciate the feedback, I don’t feel personally touched”. Contributors described wanting to know more about why they were being thanked and what users were using their packages for (e.g. “more than the one sentence of thanks, what was useful for them that they were thanking for?” (U19) and “having a rough idea about types of projects my features are used for is quite beneficial.” (C23)). Most contributors were surprised to receive the notification (Figure 5 (B)), however, few contributors who belonged to large projects were not. Most contributors felt no or low levels of awkwardness being appreciated (Figure 5 (C)). Some participants felt awkward because they did not feel like they had contributed significantly to the project for which they were being thanked (we discuss this further in 5.5). One contributor, who felt “extremely awkward” noted that a ‘hug’ from a stranger felt awkward (quote in Figure 5).
@@ 5.2. Hug Reports encouraged appreciation that was still meaningful to users and contributors (RQ1)
5.2.1. The button served as an ambient reminder of contributors’ effort (RQ1-U2)
@@
5.2.2. Low procedural effort encouraged appreciation (RQ1-U1, RQ1-C)
@@
“It felt like a nice opportunity for me that thank them without having to find out how to contact them. Right now, you don’t have any way of sending thanks and so, if you lower the barrier of sending thanks compared to ‘I have to find where to GitHub repo is, and leave an issue’, I would assume it would increase the number of people who actually send thanks.” (U2)
@@
Further, users liked having the option to express appreciation as a thanks since it was low effort enough that it didn’t disrupt their flow and it gave them a way to express appreciation when they didn’t “know what to say” (U9). Contributors acknowledged this sentiment behind the thanks: “I think it’s helpful. Because sometimes you don’t know what to say in a comment. You don’t have something particularly meaningful, but you still appreciate it” (C12). 14 of the contributors who responded to the survey had received only a thanks without any personal notes, 5 of whom reported feeling no different than normal, 3 reported feeling much more positive than normal, and the remaining 6 reported positive emotions in between the two. None of them reported feeling negatively. Since the extension gave users flexibility in how much effort they could convey, C13 suggested: “in this hierarchy of showing support you go all the way from not doing anything to supporting monetarily right? And somewhere along the rung there is this, the starring of the GitHub Repo. And at a slightly higher, more personalized level, there is thanking”.
5.3. Moments in which users expressed appreciation (RQ2-U)
Users reflected on temporal rhythms in their feelings and expression of appreciation which provide a deeper understanding of usage trends.
5.3.1. Users engaged with the extension when they “came up for air” (RQ2-U)
@@
Users described how they sent thanks in moments of transition between tasks, or in moments of “rest” (U7) between tasks. These moments often coincided with transitions between files and so, import statements at the top, where the file would open, were often the site where users sent thanks. U6 described these as moments when she was “coming up for air”, elaborating:
@@
I think [sending thanks] is something that I did when I wasn’t really in the middle of any sort of programming task. Having them at the top, like with all the imports, is kind of nice, because it’s naturally, where you kind of start within a file. So, if I open a file I’m like, ‘oh, yeah there’s all these imports like I should thank them’. It’s like, I’m already kind of in a paused state. I’m getting ready to do something, or I just finish something like.
@@
When users were focused on a task, “the icon became sort of like a background noise” (U1). U2 and U14 mentioned a shared a similar reflection: “When I’m down in a file I really wanna focus on writing the code rather than stopping and saying thanks” (U2) and “I didn’t really want to break my coding flow, so I would usually club them once I had some part of some module running” (U14).
@@
5.3.2. Users thanked packages retrospectively, repeating if they discovered new use cases (RQ2-U)
@@
Even though thanks was expressed at the top of the file, and in transitional moments, it was not expressed preemptively. Users mentioned how they felt appreciative, and thanked packages, once they had got them to work for their specific use case: “It’s like it did the thing. It has shown me that it can do the thing. And this is a good thing”(U9). As U15 put it: “I wasn’t saying good idea, but good implementation”. Similarly, U3 mentioned: “After I ran the code, then I would scroll up and say thank you.”. U14 also described scrolling up retrospectively: “whenever I saw that I had completed one module, I would just go at the top and see what modules I’ve used”. Further, users often described feeling appreciative when they discovered new functionalities. U7 mentioned: “I’d like to say thanks if I find something useful or something I didn’t know before. I think saying thanks sort of strengthens the my memory of my experience with that specific function.”. Users also mentioned how they would be inclined to thank packages again if they found new use cases, suggesting that thanks can be meaningfully different than one-time exchanges like starring on GitHub. U9 suggested: “I think if there is a unique functionality that hadn’t been captured before, so like you know, I have use case A, and then I find out there’s a use case B, I would definitely thank again.”
5.3.3. Broader rhythms in work influenced appreciation (RQ2-U)
@@
Users described how their engagement with the extension was also influenced by broader rhythms in their work, where they found themselves more prone to reflection during certain periods: “I have my Tuesday meetings. I would just link this to my meeting, and I was like after the meeting. I’ll do this. And so I would do this like bout of thanksgiving(?)” (U10). U1 described how project cycles could also prompt reflection: “if I’m at the end of a project cycle, then I might feel gratitude again”. U3 drew an analogy to borrowing a tool: “If someone gave you a tool for you to use, I feel like you would say thank you to them like once, when you get the tool and a second time when you return the tool.”
5.4. Two meanings of appreciation (RQ2-U, RQ2-C)
@@
Users and contributors interpreted appreciation as both: (1) a measure of utility, where the volume of thanks, and what is was directed at, were interpreted as an signal of the software’s utility, and (2) as an act of expressive communication that intended to convey a user’s gratitude. Users’ interactions and contributors’ interpretations depended on which of the two meanings they prioritized in a given context
5.4.1. Appreciation as a measure of utility (RQ2-U, RQ2-C)
@@
Despite our small scale deployment, users and contributors saw value in the quantitative metrics that could be derived from the thanks. Some contributors envisioned metrics as being useful for providing external evidence of their own impact (e.g. “I wish this was integrated with GitHub/LinkedIn (badge, )etc. It would be good and motivational to share it with my professional network” (C1)) as well as the project’s impact: “We would find it useful because funding agencies want to know that…It’s only the funding people who are like ‘but are people using this?”’ (C8). To access its informational value, contributors preferred a publicly shareable aggregation of data (C19, C13). C13 suggested how the system could “keep aggregating the stats on a website”.
@@
“In npm you’ve got those traffic stats like this got downloaded 500,000 times right. But the reason a lot of packages get downloaded very often is because they are part of this massive npm package that has 700 dependencies and you need to install all of them to just use like 4 lines of code from the top level directory but you never actually hit that line of code. So you don’t end up using it, so to speak. this could be a slightly more human-centric way of understanding that” (C27)
@@
5.4.2. Appreciation as an act of expressive communication (RQ2-U, RQ2-C)

### v2

@@ 4.3.3. Procedure and Participants
To understand contributors’ reactions to the notifications, the email also included an invitation to participate in our study with a link to a survey. The survey included questions on demographics including age, gender, programming proficiency, and their tenure in the project for which they were thanked. It then asked contributors how they felt on three single-item scales adapted from prior work (Kumar and Epley, 2018): (1) a scale ranging from -5 (much more negative than normal) to 5 (much more positive than normal), with the midpoint of 0 labeled no different than normal; (2) a scale ranging from -5 (not at all surprised) to 5 (extremely surprised); and (3) a scale ranging from -5 (not at all awkward) to 5 (extremely awkward). We also included an open-ended question asking contributors to further describe how the notification made them feel. The survey included three more open-ended questions: (1) Prior to this, how often did you receive messages of thanks from users?; (2) What else would you have liked to know about the senders of these thanks?; and (3) Do you have any feedback for us that you would like to share? Finally, the survey invited contributors to participate in an interview. There was no compensation for participating in the survey and all questions on the survey were indicated as optional.
@@ 4.4. Analysis
We used the usage data collected from the extension deployment as one source of data to address RQ2. From this, we created summary statistics of the interactions that took place. We also used an affinity diagramming approach to group personal notes into categories, based on their content. We used contributors’ responses to the survey, as one source of data to address RQ1 and RQ2. We created summary statistics for responses to quantitative questions and used an affinity diagramming approach to analyze the responses to open-ended questions. In this way, we used these two sources of data to derive descriptive summaries of the behaviors of users and reactions of contributors.
We used the interviews to address RQ1, RQ2, and RQ3. To do this, we conducted a reflexive thematic analysis (Braun and Clarke, 2006). Two of the authors independently performed a line-by-line open coding of transcripts from the first 7 (out of the 14) user interviews and the first 5 (out of the 10) contributor interviews. Codes generated in this phase were in part inductive, driven by the data, and in part guided by our original research questions– we remained open to capturing observations that emerged through the data while also looking out for observations that related to our main guiding questions. All authors met to discuss the analysis and iteratively refined the codes, following which the first author applied the refined set of codes to the remaining transcripts looking at whether participant experiences fit into our existing categories. Finally, all authors discussed the analysis to iteratively refine and solidify the themes, and group similar themes together. Themes were generated at a semantic level, reflecting what participants explicitly said (Braun and Clarke, 2006).
@@ 5. Findings
In this section, we begin with findings from the descriptive analysis of usage data and contributors’ responses to the questionnaire and survey (5.1). In it, we provide a descriptive summary of the extension’s use and contributors’ reactions, addressing RQ1, and RQ2. Across the next subsections, we present the main themes from our analysis of the interviews. In section 5.2, we first discuss the extent to which key decisions of the extension encouraged meaningful appreciation (RQ1). In sections 5.3, 5.4, and 5.5 we present findings that speak to RQ2, describing how appreciation was felt, expressed, received, and interpreted. Section 5.3 describes the moments in which users expressed appreciation. Section 5.4 describes the two meanings that appreciation took on: (1) as a measure of utility, where the volume of thanks, and what it was directed at, were interpreted as a signal of the software’s utility, and (2) as an act of expressive communication that intended to convey a user’s gratitude. Users’ interactions and contributors’ interpretations depended on which of the two meanings they prioritized in a given context (5.4). In section 5.5, we discuss how contributors’ reactions were influenced by how much they felt they had contributed to the object that was being appreciated. Finally, we present participants’ ideas (5.6) for further supporting exchanges of appreciation, addressing RQ3.
@@ 5.1.1. Patterns of use (RQ2)
Figure 5. Summary of contributors’ responses to the survey. We include some contributors’ responses to the open-ended questions in the survey to provide further context for some responses. (A) Most contributors reacted positively to the notification, and only 1 participant reacted negatively (whose survey response suggests it was because they found the thanks and personal note too generic). (B) Most contributors were surprised to receive the notification, however, few contributors who belonged to large projects were not. (C) Most contributors felt no or low levels of awkwardness being appreciated. One contributor, who felt ‘extremely awkward’ noted that receiving a ‘hug’ from a stranger felt awkward.
5.1.2. Contributors’ responses (RQ1, RQ2)
@@
Overall, contributors reacted positively to receiving appreciation. Contributors hoped to know more about why they were being thanked and would have liked more personalized messages. Some contributors didn’t feel like they deserved the appreciation if they hadn’t substantially contributed to the package for which they were thanked. Figure 5 summarizes the responses we received to the survey, from 26 contributors. 14 contributors reported that receiving the notification made them feel more positive than normal (Figure 5 (A)), 11 reported feeling no different than normal, and only 1 participant reacted negatively (whose survey response suggests it was because they found the thanks and personal note too generic). Responding to how often they currently received appreciation, most contributors reported it was rare, some suggesting they received “a few per year” (C15), such as in “GitHub Issues once every couple months” (C13) or “once a feature request is closed, primarily with emojis, and it is extremely rare to get a text message” (C23). Participants further described how the notification made them feel. One mentioned how “receiving the hug report was a wonderful surprise, and brightened [their] day” (C3), while another mentioned they “felt happy and feel motivated to do more” (C1). One contributor explicitly mentioned how current channels for appreciation are few: “receiving thanks for open-source work felt nice and very new since I experienced it very rarely so far. Part of why it is so rare may be that there are no convenient and established ways to do so” (C20). However, few contributors noted how there was “nothing personalized about the report” (C9). Without a more specific message, C18 mentioned: “although I do appreciate the feedback, I don’t feel personally touched”. Contributors described wanting to know more about why they were being thanked and what users were using their packages for (e.g. “more than the one sentence of thanks, what was useful for them that they were thanking for?” (U19) and “having a rough idea about types of projects my features are used for is quite beneficial.” (C23)). Most contributors were surprised to receive the notification (Figure 5 (B)), however, few contributors who belonged to large projects were not. Most contributors felt no or low levels of awkwardness being appreciated (Figure 5 (C)). Some participants felt awkward because they did not feel they had contributed significantly to the project for which they were being thanked (we discuss this further in 5.5). One contributor, who felt “extremely awkward” noted that a ‘hug’ from a stranger felt awkward (quote in Figure 5).
@@ 5.2. Hug Reports encouraged appreciation that was still meaningful to users and contributors (RQ1)
5.2.1. The button served as an ambient reminder of contributors’ effort
@@
5.2.2. Low procedural effort encouraged appreciation
@@
“It felt like a nice opportunity for me to thank them without having to find out how to contact them. Right now, you don’t have any way of sending thanks and so, if you lower the barrier of sending thanks compared to ‘I have to find where to GitHub repo is, and leave an issue’, I would assume it would increase the number of people who actually send thanks.” (U2)
@@
Further, users liked having the option to express appreciation as a thanks since it was low effort enough that it didn’t disrupt their flow and it gave them a way to express appreciation when they didn’t “know what to say” (U9). Contributors acknowledged this sentiment behind the thanks: “I think it’s helpful. Because sometimes you don’t know what to say in a comment. You don’t have something particularly meaningful, but you still appreciate it” (C12). 14 of the contributors who responded to the survey had received only a thanks without any personal notes, 5 of whom reported feeling no different than normal, 3 reported feeling much more positive than normal, and the remaining 6 reported positive emotions in between the two. None of them reported feeling negatively. Since the extension gave users flexibility in how much effort they could convey, C13 suggested: “in this hierarchy of showing support you go all the way from not doing anything to supporting monetarily right? And somewhere along the rung, there is this, the starring of the GitHub Repo. And at a slightly higher, more personalized level, there is thanking”.
5.3. Moments in which users expressed appreciation (RQ2)
5.3.1. Users engaged with the extension when they “came up for air”
@@
Users described how they sent thanks in moments of transition between tasks, or moments of “rest” (U7) between tasks. These moments often coincided with transitions between files and so, import statements at the top, where the file would open, were often the site where users sent thanks. U6 described these as moments when she was “coming up for air”, elaborating:
@@
I think [sending thanks] is something that I did when I wasn’t really in the middle of any sort of programming task. Having them at the top, like with all the imports, is kind of nice, because it’s naturally, where you kind of start within a file. So, if I open a file I’m like, ‘oh, yeah there’s all these imports like I should thank them’. It’s like, I’m already kind of in a paused state. I’m getting ready to do something, or I just finish something.
@@
When users were focused on a task, “the icon became sort of like a background noise” (U1). U2 and U14 shared a similar reflection: “When I’m down in a file I really wanna focus on writing the code rather than stopping and saying thanks” (U2) and “I didn’t really want to break my coding flow, so I would usually club them once I had some part of some module running” (U14).
@@
5.3.2. Users thanked packages retrospectively, repeating if they discovered new use cases
@@
Even though thanks was expressed at the top of the file, and in transitional moments, it was not expressed preemptively. Users mentioned how they felt appreciative, and thanked packages, once they had got them to work for their specific use case: “It’s like it did the thing. It has shown me that it can do the thing. And this is a good thing”(U9). As U15 put it: “I wasn’t saying good idea, but good implementation”. Similarly, U3 mentioned: “After I ran the code, then I would scroll up and say thank you.”. U14 also described scrolling up retrospectively: “whenever I saw that I had completed one module, I would just go at the top and see what modules I’ve used”. Further, users often described feeling appreciative when they discovered new functionalities. U7 mentioned: “I’d like to say thanks if I find something useful or something I didn’t know before. I think saying thanks sort of strengthens my memory of my experience with that specific function.”. Users also mentioned how they would be inclined to thank packages again if they found new use cases, suggesting that thanks can be meaningfully different than one-time exchanges like starring on GitHub. U9 suggested: “I think if there is a unique functionality that hadn’t been captured before, so like you know, I have use case A, and then I find out there’s a use case B, I would definitely thank again.”
5.3.3. Broader rhythms in work influenced appreciation
@@
Users described how their engagement with the extension was also influenced by broader rhythms in their work, where they found themselves more prone to reflection during certain periods: “I have my Tuesday meetings. I would just link this to my meeting, and I was like after the meeting. I’ll do this. And so I would do this like bout of thanksgiving(?)” (U10). U1 described how project cycles could also prompt reflection: “if I’m at the end of a project cycle, then I might feel gratitude again”. U3 drew an analogy to borrowing a tool: “If someone gave you a tool for you to use, I feel like you would say thank you to them once when you get the tool and a second time when you return the tool.”
5.4. Two meanings of appreciation (RQ2)
@@
Users and contributors interpreted appreciation as both: (1) a measure of utility, where the volume of thanks, and what it was directed at, were interpreted as a signal of the software’s utility, and (2) as an act of expressive communication that intended to convey a user’s gratitude. Users’ interactions and contributors’ interpretations depended on which of the two meanings they prioritized in a given context.
5.4.1. Appreciation as a measure of utility
@@
Despite our small-scale deployment, users and contributors saw value in the quantitative metrics that could be derived from the thanks. Some contributors envisioned metrics as being useful for providing external evidence of their own impact (e.g. “I wish this was integrated with GitHub or LinkedIn badges. It would be good and motivational to share it with my professional network” (C1)) as well as the project’s impact: “We would find it useful because funding agencies want to know that…It’s only the funding people who are like ‘But are people using this?”’ (C8). To access its informational value, contributors preferred a publicly shareable aggregation of data (C19, C13). C13 suggested how the system could “keep aggregating the stats on a website”.
@@
“In npm you’ve got those traffic stats like this got downloaded 500,000 times right. But the reason a lot of packages get downloaded very often is because they are part of this massive npm package that has 700 dependencies and you need to install all of them to just use like 4 lines of code from the top-level directory but you never actually hit that line of code. So you don’t end up using it, so to speak. this could be a slightly more human-centric way of understanding that” (C27)
@@
5.4.2. Appreciation as an act of expressive communication